OpenStudy (anonymous):

The vapour density (hydrogen=1) of a mixture consisting of NO2 and N2O4 is 38.3 at 26.7 degree Celcius. Calucute the number of moles of NO2in 100 g of the mixture. Please Explain!!!

OpenStudy (aaronq):

density is defined as, d=m/V 100g/(38.3g/mL)=2.61096 mL seems like very little, but since you didn't provide units, i can only speculate PV=nRT (1) P(0.00261L)=n(0.08206)(299.85) ma= mass of NO2 mb= mass of N2O2 since, 100g=ma+mb -> ma=100g-mb na= moles of NO2; nb= moles of N2O2 Ma=molar mass of NO2; Mb=molar mass of N2O2 (2) n(T)= na+nb = ma/Ma + mb/Mb = (100g-mb)/Ma + mb/Mb P(0.00261L)=n(0.08206)(299.85) ->P/n=constant grouping values as 1 constant: \[\frac{ P }{\frac{ (100g-m _{N _{2} O _{2}}) }{ 46g/mol } +\frac{ m _{N _{2}O _{2}} }{ 60g/mol }}=9427.46 \frac{ atm }{ mol }\] thats about where i get stuck, it seems like you would another value, like pressure, mole fraction or something.

OpenStudy (anonymous):

The answer is 0.437 mole!